OpenStudy (anonymous):

Find the Volume of a cylinder with a radius 11 cm and a height of 7 cm.

OpenStudy (anonymous):

so I did

OpenStudy (anonymous):

V=pi*11^2*7 V=pi*121*7 V=pi*847 v=2,659.58 did I solve right?

OpenStudy (ujjwal):

yeah you did it right.. well the unit will be cm^3

OpenStudy (anonymous):

ok thanks :)